### v1

@@ Abstract
Deep neural networks (DNNs) are known to be susceptible to adversarial examples, leading to significant performance degradation. In black-box attack scenarios, a considerable attack performance gap between the surrogate model and the target model persists. This work focuses on enhancing the transferability of adversarial examples to narrow this performance gap. We observe that the gradient information around the clean image, i.e. Neighbourhood Gradient Information, can offer high transferability. Leveraging this, we propose the NGI-Attack, which incorporates Example Backtracking and Multiplex Mask strategies, to use this gradient information and enhance transferability fully. Specifically, we first adopt Example Backtracking to accumulate Neighbourhood Gradient Information as the initial momentum term. Multiplex Mask, which forms a multi-way attack strategy, aims to force the network to focus on non-discriminative regions, which can obtain richer gradient information during only a few iterations. Extensive experiments demonstrate that our approach significantly enhances the adversarial transferability. Especially, when attacking numerous defense models, we achieve an average attack success rate of 95.8%. Notably, our method can plugin with any off-the-shelf algorithm to improve their attack performance without additional time cost.
@@ 1.  Introduction
Deep neural networks (DNNs) have made significant progress and are widely applied in security-sensitive applications, such as autonomous driving (Yang et al., 2023; Hong et al., 2023), face anti-spoofing (Zheng et al., 2024b, a), and face recognition (Wang et al., 2018). However, recent work (Goodfellow et al., 2015; Madry et al., 2018; Chen et al., 2022a, b; Huang et al., 2022; Chen et al., 2023a; Jiang et al., 2023) has revealed that DNNs are highly susceptible and vulnerable to adversarial examples, which introduce imperceptible perturbations to the input and ultimately cause the model to produce erroneous predictions.
@@
Adversarial attacks are typically categorized into white-box and black-box settings. In the white-box settings, gradient-based attacks such as Projected Gradient Descent (PGD) (Madry et al., 2018) exhibit remarkably high attack success rates on targeted models, capable of accessing all information (e.g., architecture or gradients) of the targeted model. However, in real-world scenarios, attackers often need to generate adversarial examples under black-box settings, as they are unaware of any information of the target model (Cheng et al., 2019, 2020; Li et al., 2020). From this perspective, adversarial transferability is crucial, enabling adversaries to exploit surrogate models to generate adversarial examples and transfer them to target models to perform attacks. Therefore, the generation of highly transferable adversarial examples (Dong et al., 2018; Xie et al., 2019; Lin et al., 2020; Wang et al., 2021; Dong et al., 2019; Qin et al., 2022; Chen et al., 2023b) has gradually become a hot topic in the community of adversarial examples.
@@
Lin et al. (2020) regard the generation of adversarial examples as an optimization process, and the white-box attack induces poor transferability by causing overfitting of the targeted model. Numerous works have been proposed to alleviate overfitting by enhancing adversarial transferability, such as advanced optimization (Dong et al., 2018; Lin et al., 2020, 2019; Wang & He, 2021), input transformation (Xie et al., 2019; Wang et al., 2021; Dong et al., 2019; Lin et al., 2019), and model augmentation (Long et al., 2022). However, a significant disparity in attack performance persists between the transfer and the optimal white-box settings, necessitating further endeavors to enhance transferability.
@@
Upon recalling these methods, advanced optimization is the most fundamental strategy, e.g. MI-FGSM (Dong et al., 2018), which accumulates gradients to become momentum and leverages momentum to aid the optimization of adversarial examples towards escaping local optima. Numerous optimization-based methods have concentrated on refining gradient accumulation and momentum application (Lin et al., 2019; Wang & He, 2021; Qin et al., 2022), aiming to fully exploit the capabilities of MI-FGSM. These enhancements aim to identify optimal strategies for gradient information and enhance adversarial transferability. However, they exhibit limited transferability because they do not distinguish the transferability differences of gradient information across iterations, instead applying it uniformly in each iteration. Another strategy, input transformation-based methods, seeks to mitigate attack overfitting by boosting the transferability of adversarial examples through various input augmentation strategies. While combining gradient optimization and input transformation approaches has been shown to significantly increase transferability (Lin et al., 2019), there remains a notable gap in achieving satisfactory performance, especially when attacking numerous defense models.
@@
To tackle the issue above, we propose a novel attack based on Neighbourhood Gradient Information, termed NGI-Attack, incorporating two key strategies: Example Backtracking and Multiplex Mask. From the perspective of advanced optimization, we observe that the gradient information generated from the initial iteration, when the adversarial examples are still around the clean images, can offer high transferability. We introduce this novel concept as Neighbourhood Gradient Information. To demonstrate its effectiveness, we perform a preliminary analysis by incorporating additional iterations before regular attacks to accumulate this specific gradient information via momentum, as detailed in Section 3.2. Utilizing the Example Backtracking method, we collect this gradient information directly. However, when the number of iterations for the attack is fixed, allocating some of these iterations to collect Neighbourhood Gradient Information consequently necessitates a larger step size to ensure the attack reaches its maximum perturbation magnitude. Nonetheless, simply increasing the step size results in poor transferability on adversarially trained models. To address this, we introduce an input transformation approach, specifically the Multiplex Mask strategy, which forms a multi-way attack. This strategy employs two distinct branches: one applies mask-enhanced images to direct the network’s focus to non-discriminative regions, while the other employs clean images. By incorporating two pathways, we can obtain richer gradient information and reach enough perturbation magnitude. Furthermore, our method is plug-and-play and requires no additional time overhead, making it easy to integrate with existing strategies. Our contributions are followed as:
@@
  • Extensive experiments show that our method achieves state-of-the-art performance. In particular, we outperform existing methods by margins of 2.6% to 19.3% under normally trained models and 2.4% to 13.5% under adversarially trained models. Furthermore, we attain an average success rate of 95.8% against defense models.
@@ 2.  Related Work
Since adversarial examples are found to exploit DNNs vulnerabilities (Szegedy et al., 2014), many adversarial attack methods have been proposed, including white-box and black-box attacks. Unlike white-box attacks, black-box attacks cannot directly access the model’s structure or gradient information. Adversarial examples under black-box attacks are mainly generated via (1) Query-based attacks using iterative queries to the target model. (2) Transfer-based attacks applying a surrogate model and then transferring to the target models to perform the attack. In this work, we focus on enhancing the transferability of adversarial examples under black-box attack. A general introduction to transfer attacks is provided as follows. Concurrently, we furnish an overview of recent advancements in adversarial defense.
@@
Fast Gradient Sign Method (FGSM). FGSM (Goodfellow et al., 2015), the pioneering algorithm that uses gradient optimization techniques, generates adversarial examples by maximizing the loss function in a single update step:
$x^{adv}=x+\epsilon\cdot\text{sign}\left(\nabla_{x}J(x,y;\theta)\right),$
@@
Iterative Fast Gradient Sign Method (I-FGSM). I-FGSM (Kurakin et al., 2018) introduces a small step size $\alpha$, extending FGSM to an iterative version, thereby improving white-box attack performance:
$x_{0}^{adv}=x,$
$x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)).$
Momentum Iterative Fast Gradient Sign Method (MI-FGSM). To escape the suboptimal local solution of the surrogate model and to boost the effect of black-box attacks, Dong et al. (2017) introduce the momentum term and integrate it into I-FGSM to improve transferability:
$g^{t+1}=\mu\cdot g^{t}+\frac{\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)}{\|\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)\|_{1}},$
$x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(g^{t+1}),$
@@
Nesterov Iterative Fast Gradient Sign Method (NI-FGSM). Lin et al. (2019) employ the Nesterov accelerated gradient to implement a lookahead in each iteration, thereby circumventing local optimal solutions:
$x_{t}^{nes}=x_{t}^{adv}+\alpha\cdot\mu\cdot g_{t},$
$g^{t+1}=\mu\cdot g^{t}+\frac{\nabla_{x_{t}^{nes}}J(x_{t}^{nes},y;\theta)}{\|\nabla_{x_{t}^{nes}}J(x_{t}^{nes},y;\theta)\|_{1}},$
$x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(g^{t+1}).$
Variance Tuning Gradient-based Attacks (VMI-FGSM). In addition,Wang & He (2021) considers the gradient variance, adjusting the current gradient by its variance to ensure stable updates:
$g^{t+1}=\mu\cdot g^{t}+\frac{\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)+v_{t}}{\|\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)+v_{t}\|_{1}},$
$v_{t+1}=V(x_{t}^{adv}),$
$V(x)=\frac{1}{N}\sum_{i=1}^{N}\left(\nabla_{x_{i}}J(x_{i},y;\theta)-\nabla_{x}J(x,y;\theta)\right).$
Reverse Adversarial Perturbation (RAP). Unlike traditional approaches that focus on a single adversarial point’s loss, RAP (Qin et al., 2022) aims to locate an adversarial example in a region of consistently low loss. Nonetheless, RAP necessitates a greater number of iterative steps to achieve optimal transferability.
@@
Diverse Input Method (DI-FGSM). DI-FGSM (Xie et al., 2019) stochastically resizes and pads the input image with a specified probability. The stochastic transformation function $T(X_{t}^{adv};p)$ is
$T(X_{t}^{adv};p)=\begin{cases}T(X_{t}^{adv})&\text{with probability }p,\\ X_{t}^{adv}&\text{with probability }1-p.\end{cases}$
Translation-Invariant Method (TI-FGSM). TI-FGSM (Dong et al., 2019) improves transferability by applying convolutional transformations to gradients with a Gaussian kernel, approximating gradients across an image collection.
$x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(W*\nabla_{x}J(x_{t}^{adv},y)).$
@@
Scale-Invariant Method (SI-FGSM). SI-FGSM (Lin et al., 2019) leverages the scaling invariance of images by calculating gradients across various scaling factors.
Admix. Admix (Wang et al., 2021) further enhances input images by adding a portion of other images, retaining original labels, and computing gradients of the mixed images:
$\tilde{x}=\gamma\cdot(x+\eta\cdot x^{\prime}),$
$\overline{g}_{t+1}=\frac{1}{m_{1}\cdot m_{2}}\sum_{x^{\prime}\in X^{\prime}}\sum_{i=0}^{m_{1}-1}\nabla_{x_{t}^{adv}}J(\gamma_{i}\cdot(x_{t}^{adv}+\eta\cdot x^{\prime}),y;\theta),$
$g_{t+1}=\mu\cdot g_{t}+\frac{\overline{g}_{t+1}}{\|\overline{g}_{t+1}\|_{1}},$
@@
SIA. SIA (Wang et al., 2023) enhances transferability by applying diverse input transformations to localized image blocks for gradient computation.
BSR. Wang et al. (2024) introduce Block Shuffle and Rotation (BSR) to enhance adversarial transferability. BSR divides the input image into multiple blocks, then randomly shuffles and rotates these blocks to disrupt the intrinsic relationships within the image, demonstrating superior performance in black-box attack settings.
@@
Spectral Simulation Attack (SSA). SSA (Long et al., 2022) concentrates on the frequency domain, employing spectral transformations on the inputs to simulate model enhancement methods, thereby boosting the transferability of adversarial examples. However, this method requires a considerable amount of time.
@@ 2.4. Adversarial Defense
In this section, we offer a brief introduction to adversarial defenses. Various defense methods have been proposed to mitigate the threat posed by adversarial examples. Adversarial training (Tramèr et al., 2018), which incorporates adversarial examples into training data, is widely considered one of the most effective approaches to enhance model robustness. However, these methods often entail a significant computational overhead. Another category of methods primarily focuses on efficiently purifying adversarial examples. Specifically, Liao et al. (2018) propose a High-level Representation Guided Denoiser (HGD). Xie et al. (2017) introduce random resizing and padding on the input (R&P). Feature Distillation (FD) (Liu et al., 2019) put forward a JPEG-based defensive compression framework. Comdefend  (Jia et al., 2019) suggest an end-to-end compression model. Cohen et al. (2019) employ a technique known as Randomized Smoothing (RS). Meanwhile, Xu et al. (2018) introduce bit reduction (Bit-Red) for the detection of adversarial examples. The Neural Representation Purifier (NRP) model by Naseer et al. (2020) employs a self-supervised adversarial training approach in the input space, focusing on cleansing adversarially perturbed images.

### v2

@@ Abstract
Deep neural networks (DNNs) are known to be susceptible to adversarial examples, leading to significant performance degradation. In black-box attack scenarios, a considerable attack performance gap between the surrogate model and the target model persists. This work focuses on enhancing the transferability of adversarial examples to narrow this performance gap. We observe that the gradient information around the clean image, i.e., Neighbourhood Gradient Information (NGI), can offer high transferability. Based on this insight, we introduce NGI-Attack, incorporating Example Backtracking and Multiplex Mask strategies to exploit this gradient information and enhance transferability. Specifically, we first adopt Example Backtracking to accumulate Neighbourhood Gradient Information as the initial momentum term. Then, we utilize Multiplex Mask to form a multi-way attack strategy that forces the network to focus on non-discriminative regions, which can obtain richer gradient information during only a few iterations. Extensive experiments demonstrate that our approach significantly enhances adversarial transferability. Especially, when attacking numerous defense models, we achieve an average attack success rate of 95.2%. Notably, our method can seamlessly integrate with any off-the-shelf algorithm, enhancing their attack performance without incurring extra time costs.
@@ 1. Introduction
Deep neural networks (DNNs) have made significant progress (Arash1; Arash2; Arash3; Arash4; Arash5; Arash6) and are widely applied in security-sensitive applications, such as autonomous driving (yang2023aide; hong2023lvos; HOANG2024104031), face anti-spoofing (10448479; ZhengLWWMLDW24), and face recognition (wang2018cosface). However, recent work (fgsm; pgd; chen2022shape; chen2022towards; huang2022cmua; chen2023devopatch; jiang2023efficient; Khaleel_Habeeb_Alnabulsi_2024; hybridattack) has revealed that DNNs are highly susceptible and vulnerable to adversarial examples, which introduce imperceptible perturbations to the input and ultimately cause the model to produce erroneous predictions.
@@
Adversarial attacks are typically categorized into white-box and black-box settings. In the white-box settings, gradient-based attacks such as Projected Gradient Descent (PGD) (pgd) exhibit remarkably high attack success rates on targeted models, capable of accessing all information (e.g., architecture or gradients) of the targeted model. However, in real-world scenarios, attackers often need to generate adversarial examples under black-box settings, as they are unaware of any information of the target model (optattack; sign-opt; QEBA). From this perspective, adversarial transferability is crucial, enabling adversaries to exploit surrogate models to generate adversarial examples and transfer them to target models to perform attacks. Therefore, the generation of highly transferable adversarial examples (mi; xie2019improving; sini; wang2021admix; dong2019evading; NEURIPS2022_c0f9419c; chen2023aca) has gradually become a hot topic in the community of adversarial examples.
@@
Lin et al. sini regard the generation of adversarial examples as an optimization process, and the white-box attack induces poor transferability by causing overfitting of the targeted model. Numerous works have been proposed to alleviate overfitting by enhancing adversarial transferability, such as advanced optimization (mi; sini; DBLP:journals/corr/abs-1908-06281; DBLP:conf/cvpr/Wang021), input transformation (xie2019improving; wang2021admix; dong2019evading; DBLP:journals/corr/abs-1908-06281; SIA; wang2024boosting), and model augmentation (ssa). However, a significant disparity in attack performance persists between the transfer and the optimal white-box settings, necessitating further endeavors to enhance transferability.
@@
Upon recalling these methods, advanced optimization is the most fundamental strategy, e.g. MI-FGSM (mi), which accumulates gradients to become momentum and leverages momentum to aid the optimization of adversarial examples towards escaping local optima. Numerous optimization-based methods have concentrated on refining gradient accumulation and momentum application (DBLP:journals/corr/abs-1908-06281; DBLP:conf/cvpr/Wang021; NEURIPS2022_c0f9419c), aiming to fully exploit the capabilities of MI-FGSM. These enhancements aim to identify optimal strategies for gradient information and enhance adversarial transferability. However, they exhibit limited transferability because they do not distinguish the transferability differences of gradient information across iterations, instead applying it uniformly in each iteration. Another strategy, input transformation-based methods, seeks to mitigate attack overfitting by boosting the transferability of adversarial examples through various input augmentation strategies. While combining gradient optimization and input transformation approaches has been shown to significantly increase transferability (DBLP:journals/corr/abs-1908-06281), there remains a notable gap in achieving satisfactory performance, especially when attacking numerous defense models.
@@
To tackle the issue above, we propose a novel attack based on Neighbourhood Gradient Information, termed NGI-Attack, incorporating two key strategies: Example Backtracking and Multiplex Mask. From the perspective of advanced optimization, we observe that the gradient information generated from the initial iteration, when the adversarial examples are still around the clean images, can offer high transferability. We introduce this novel concept as Neighbourhood Gradient Information (NGI). To demonstrate its effectiveness, we perform a preliminary analysis by incorporating additional iterations before regular attacks to accumulate this specific gradient information via momentum, as detailed in Section 3.2. Utilizing the Example Backtracking method, we collect this gradient information directly. However, when the number of iterations for the attack is fixed, allocating some of these iterations to collect Neighbourhood Gradient Information consequently necessitates a larger step size to ensure the attack reaches its maximum perturbation magnitude. Nonetheless, simply increasing the step size results in poor transferability on adversarially trained models. To address this, we introduce an input transformation approach, specifically the Multiplex Mask strategy, which forms a multi-way attack. This strategy employs two distinct branches: one applies MaskProcess images to direct the network’s focus to non-discriminative regions, while the other employs clean images. By incorporating two pathways, we can obtain richer gradient information and reach enough perturbation magnitude. Furthermore, our method is plug-and-play and requires no additional time overhead, making it easy to integrate with existing strategies. Our contributions are followed as:
@@
  • Extensive experiments show that our method achieves state-of-the-art performance. In particular, we outperform existing methods by margins of 2.6% to 19.3% under normally trained models and 2.4% to 13.5% under adversarially trained models. Furthermore, we attain an average success rate of 95.2% against defense models.
@@ 2. Related Work
Since adversarial examples are found to exploit DNNs vulnerabilities (L_BFGS), many adversarial attack methods have been proposed, including white-box and black-box attacks. Unlike white-box attacks, black-box attacks cannot directly access the model’s structure or gradient information. Adversarial examples under black-box attacks are mainly generated via (1) Query-based attacks using iterative queries to the target model. (2) Transfer-based attacks applying a surrogate model and then transferring to the target models to perform the attack. In this work, we focus on enhancing the transferability of adversarial examples under black-box attack. A general introduction to transfer attacks is provided as follows. Concurrently, we furnish an overview of recent advancements in adversarial defense.
@@
Fast Gradient Sign Method (FGSM). FGSM (fgsm), the pioneering algorithm that uses gradient optimization techniques, generates adversarial examples by maximizing the loss function in a single update step:
$x^{adv}=x+\epsilon\cdot\text{sign}\left(\nabla_{x}J(x,y;\theta)\right),$ (1)
@@
Iterative Fast Gradient Sign Method (I-FGSM). I-FGSM (ifgsm) introduces a small step size $\alpha$, extending FGSM to an iterative version, thereby improving white-box attack performance:
$x_{0}^{adv}=x,\quad x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)).$ (2)
Momentum Iterative Fast Gradient Sign Method (MI-FGSM). To escape the suboptimal local solution of the surrogate model and to boost the effect of black-box attacks, dong2017discovering introduce the momentum term and integrate it into I-FGSM to improve transferability:
$g^{t+1}=\mu\cdot g^{t}+\frac{\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)}{\|\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)\|_{1}},\quad x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(g^{t+1}),$ (3)
@@
Nesterov Iterative Fast Gradient Sign Method (NI-FGSM). DBLP:journals/corr/abs-1908-06281 employ the Nesterov accelerated gradient to implement a lookahead in each iteration, thereby circumventing local optimal solutions:
$x_{t}^{nes}=x_{t}^{adv}+\alpha\cdot\mu\cdot g_{t},\quad g^{t+1}=\mu\cdot g^{t}+\frac{\nabla_{x_{t}^{nes}}J(x_{t}^{nes},y;\theta)}{\|\nabla_{x_{t}^{nes}}J(x_{t}^{nes},y;\theta)\|_{1}},\quad x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(g^{t+1}).$ (4)
Variance Tuning Gradient-based Attacks (VMI-FGSM). In addition, DBLP:conf/cvpr/Wang021 considers the gradient variance, adjusting the current gradient by its variance to ensure stable updates:
$g^{t+1}=\mu\cdot g^{t}+\frac{\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)+v_{t}}{\|\nabla_{x_{t}^{adv}}J(x_{t}^{adv},y;\theta)+v_{t}\|_{1}},\quad v_{t+1}=V(x_{t}^{adv}),\quad V(x)=\frac{1}{N}\sum_{i=1}^{N}\left(\nabla_{x_{i}}J(x_{i},y;\theta)-\nabla_{x}J(x,y;\theta)\right).$ (5)
Reverse Adversarial Perturbation (RAP). Unlike traditional approaches that focus on a single adversarial point’s loss, RAP (NEURIPS2022_c0f9419c) aims to locate an adversarial example in a region of consistently low loss. Nonetheless, RAP necessitates a greater number of iterative steps to achieve optimal transferability.
@@
Diverse Input Method (DI-FGSM). DI-FGSM (xie2019improving) stochastically resizes and pads the input image with a specified probability. The stochastic transformation function $T(X_{t}^{adv};p)$ is
$T(X_{t}^{adv};p)=\begin{cases}T(X_{t}^{adv})&\text{with probability }p,\\ X_{t}^{adv}&\text{with probability }1-p.\end{cases}$ (6)
Translation-Invariant Method (TI-FGSM). TI-FGSM (dong2019evading) improves transferability by applying convolutional transformations to gradients with a Gaussian kernel, approximating gradients across an image collection.
$x_{t+1}^{adv}=x_{t}^{adv}+\alpha\cdot\text{sign}(W*\nabla_{x}J(x_{t}^{adv},y)).$ (7)
@@
Scale-Invariant Method (SI-FGSM). SI-FGSM (DBLP:journals/corr/abs-1908-06281) leverages the scaling invariance of images by calculating gradients across various scaling factors.
Admix. Admix (wang2021admix) further enhances input images by adding a portion of other images, retaining original labels, and computing gradients of the mixed images:
$\tilde{x}=\gamma\cdot(x+\eta\cdot x^{\prime}),\quad\overline{g}_{t+1}=\frac{1}{m_{1}\cdot m_{2}}\sum_{x^{\prime}\in X^{\prime}}\sum_{i=0}^{m_{1}-1}\nabla_{x_{t}^{adv}}J(\gamma_{i}\cdot(x_{t}^{adv}+\eta\cdot x^{\prime}),y;\theta),\quad g_{t+1}=\mu\cdot g_{t}+\frac{\overline{g}_{t+1}}{\|\overline{g}_{t+1}\|_{1}},$ (8)
@@
SIA. Structure Invariant Transformation (SIA) (SIA) enhances transferability by applying diverse input transformations to localized image blocks for gradient computation.
BSR. wang2024boosting introduce Block Shuffle and Rotation (BSR) to enhance adversarial transferability. BSR divides the input image into multiple blocks, then randomly shuffles and rotates these blocks to disrupt the intrinsic relationships within the image, demonstrating superior performance in black-box attack settings.
@@
Spectral Simulation Attack (SSA). SSA (ssa) concentrates on the frequency domain, employing spectral transformations on the inputs to simulate model enhancement methods, thereby boosting the transferability of adversarial examples. However, this method requires a considerable amount of time.
@@ 2.4. Adversarial Defense
In this section, we offer a brief introduction to adversarial defenses. Various defense methods have been proposed to mitigate the threat posed by adversarial examples. Adversarial training (Florian2018Ensemble), which incorporates adversarial examples into training data, is widely considered one of the most effective approaches to enhance model robustness. However, these methods often entail a significant computational overhead. Another category of methods primarily focuses on efficiently purifying adversarial examples. Specifically, DBLP:conf/cvpr/LiaoLDPH018 propose a High-level Representation Guided Denoiser (HGD). xie2017mitigating introduce random resizing and padding on the input (R&P). Feature Distillation (FD) (liu2019feature) put forward a JPEG-based defensive compression framework. Comdefend  (DBLP:conf/cvpr/JiaWCF19) suggest an end-to-end compression model. DBLP:conf/icml/CohenRK19 employ a technique known as Randomized Smoothing (RS). Meanwhile, DBLP:conf/ndss/Xu0Q18 introduce bit reduction (Bit-Red) for the detection of adversarial examples. The Neural Representation Purifier (NRP) model by DBLP:conf/cvpr/NaseerKHKP20 employs a self-supervised adversarial training approach in the input space, focusing on cleansing adversarially perturbed images. Recently, diffusion models have emerged as powerful generative models (ho2020denoising; song2021scorebased). Nie et al. (nie2022DiffPure) introduce DiffPure, which leverages pre-trained diffusion models for adversarial purification, and demonstrate its power.